# Spectral Small-Incremental-Entangling: Breaking Quasi-Polynomial Complexity Barriers in Long-Range Interacting Systems

### v1

@@ Abstract
How the detailed structure of quantum complexity emerges from quantum dynamics remains a fundamental challenge highlighted by advances in quantum simulators and information processing. The paradigmatic principle is the Lieb–Robinson bound, setting a speed limit on information propagation but relying on geometric locality. As a more general concept, the Small-Incremental-Entangling (SIE) theorem was proposed, providing a universal constraint on the rate of entanglement generation. While the SIE theorem limits the total amount of entanglement, it leaves a major open problem in fully characterizing the fine entanglement structure. Here we introduce the concept of Spectral-Entangling Strength, which captures the structural entangling power of an operator, and use it to establish a new Spectral SIE theorem: we derive a universal speed limit for Rényi entanglement growth at $\alpha\geq 1/2$, revealing a robust $1/s^{2}$ tail in the entanglement spectrum. Remarkably, our bound at $\alpha=1/2$ is both qualitatively and quantitatively optimal, establishing the universal threshold beyond which entanglement growth becomes unbounded. This result exposes the detailed structure of Schmidt coefficients and, in turn, enables rigorous truncation-based error control, providing a quantitative link between entanglement structure and computational complexity. Building on these results, our framework establishes a generalized entanglement area law under adiabatic paths, thus extending a central principle of quantum many-body physics to general interactions. As a more practical application, we show that one-dimensional long-range interacting systems admit polynomial bond-dimension approximations for ground states, time-evolved states, and thermal states. This closes the long-standing quasi-polynomial gap and demonstrates that such systems can be simulated with polynomial complexity comparable to short-range models. In particular, by controlling Rényi entanglement, we derive the first rigorous precision-guarantee bound for the time-dependent density-matrix-renormalization-group algorithm. Overall, our results extend the scope of the SIE theorem and establish a unified framework that reveals the detailed structure of quantum complexity.
@@ I Introduction
Another well-known information-theoretic constraint that does not rely on geometric locality is the Small-Incremental-Entangling (SIE) theorem [21]. This theorem considers a bipartition of a quantum system into subsystems $A$ and $B$, and asserts that the rate at which entanglement entropy is generated between them can be upper bounded solely in terms of the interaction strength and the Hilbert space dimensions of the directly coupled subsystems $A_{0}\subset A$ and $B_{0}\subset B$. In systems with spatial locality, the SIE bound can be derived from the Lieb–Robinson bound [14]. However, its general validity across arbitrary quantum systems was first conjectured by Kitaev and Bravyi, and later fully proven in 2013 [22, 23, 24, 25]. Because the rate of entanglement growth depends only on the interaction across the boundary between $A$ and $B$, the SIE theorem is often referred to as the Dynamical Area Law. This term emphasizes the fact that entanglement generation in time is governed by the surface—rather than the volume—of the interacting regions. The SIE theorem has served as a key methodological tool in a wide range of fundamental problems, such as supporting the area law conjecture for non-critical ground states in higher-dimensional systems [22, 26], analyzing complexity growth in quantum circuits [27, 28], and characterizing measurement-induced phase transitions [29]. These applications demonstrate the power and universality of the SIE framework. Nevertheless, as we will discuss in the next section, the entanglement generation rate alone is often insufficient to fully capture the complexity of quantum many-body dynamics.
@@
In this work, we move beyond these limitations by establishing a spectral formulation of entanglement growth. Our framework not only provides rigorous dynamical constraints on the entanglement spectrum itself, but also achieves both qualitative and quantitative optimality: at the critical threshold, we establish a universal upper bound corresponding to a $1/s^{2}$ decay of squared Schmidt coefficients, which is sharp and cannot be improved. This perspective sharpens the dynamical area law into a more fine-grained structural principle and extends its reach to broader consequences, including generalized area laws of gapped ground states, efficient matrix-product-state approximations for long-range systems, and rigorous performance guarantees for numerical algorithms such as t-DMRG (time-dependent density matrix renormalization group). Conceptually, the ability to address the entanglement spectrum directly opens the door to a more refined understanding of quantum complexity, enabling applications that go well beyond entropy-based diagnostics.
@@ II.2 Spectral Small-Incremental-Entangling (SIE)
In general, this decay law cannot be further improved without imposing additional geometric assumptions such as finite-dimensionality or short-range interactions [32, 8, 9]. The result can be derived from an upper bound on the growth rate of the $\alpha$-Rényi entanglement $E_{\alpha}$, as defined precisely in Eq. (3). In order to emphasize the spectral nature of the bound, we call this extension the Spectral Small-Incremental-Entangling (Spectral SIE) theorem.
@@
An important observation is that $\alpha=1/2$ constitutes a universal threshold. For $\alpha<1/2$, the entanglement generation rate becomes unbounded, as proven in Proposition 8. Moreover, for $\alpha=1/2$ the constant $c_{\alpha=1/2}$ is not only universal but also quantitatively optimal, as shown in Proposition 7. A significant advantage of considering the Rényi entanglement with $\alpha<1$ is that it enables rigorous error analysis in terms of Schmidt rank truncation [30, 31], thereby allowing a quantitative estimation of how the entanglement structure impacts the computational complexity of simulating quantum dynamics.
@@ II.3 Low-Schmidt-rank approximation of general operators
As discussed above, it has been established that time-evolved states admit efficient low-rank approximations. A natural question is whether a similar low-rank approximation exists at the level of the time-evolution operator itself. The problem of approximating an operator in terms of operator norm is, however, significantly more challenging. To address this, we demonstrate that there exists a complexity separation between Schmidt-rank approximability of time-evolved quantum states and that of the corresponding time-evolution unitary.
[Proposition 9 (informal)] There exist Hamiltonians $H_{AB}$ with $\bar{{\mathcal{J}}}(V_{AB})=1$ for which the unitary $e^{-iH_{AB}t}$ does not admit any low-rank approximation in a certain time regime.
@@ II.4 Generalized entanglement area law
The entanglement area law refers to the property that, for a ground state of a finite-dimensional lattice system partitioned into two regions $A$ and $B$, the entanglement entropy scales with the size of the boundary between the two regions [11, 38]. In particular, the area law conjecture argues that the area law always holds for non-critical ground states with a spectral gap, and it remains one of the central open problems in quantum many-body physics [39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. The notion of an area law can be extended to more general bipartite interaction settings [49]. Specifically, when a system is divided into $A$ and $B$, the Schmidt rank of the interaction operator between $A$ and $B$ can be interpreted as an effective boundary size, thereby allowing a generalized formulation of the area law. For example, in a four-partite setup $(A_{1},A_{0},B_{0},B_{1})$ where the interaction exists only between $A_{0}$ and $B_{0}$, the relevant boundary size is determined by the dimensions of $A_{0}$ and $B_{0}$. Typical examples include generalized area laws for local Hamiltonians defined on graphs, which have been widely discussed in the literature. According to the conventional SIE theorem [22, 23], the generalized area law can be shown to hold for entanglement dynamics even in such general setups. It is also widely known that generalized area laws for mutual information hold for quantum Gibbs states at finite temperature [50, 51]. However, whether the same statement applies to gapped ground states has long been debated, and this question was resolved in the negative by Aharonov et al. in 2014 [49]. At present, aside from special classes [52], the precise conditions under which a generalized area law holds remain an important open question.
@@
In this work, we establish that the generalized area law holds under the assumption of the existence of an adiabatic path [53]. Given a parameter-dependent Hamiltonian $H(s)$, we define an adiabatic path as a continuous trajectory in parameter space that connects a reference Hamiltonian with a known area-law-satisfying ground state to a target Hamiltonian, while maintaining a finite spectral gap throughout. This condition is standard in proofs of the area law for spatially local systems [26, 22, 54], where it is combined with the Lieb–Robinson bound, and where the SIE theorem provides a key estimate on entanglement growth along the adiabatic evolution. In more general settings, such as systems defined on infinite-dimensional graphs or general two-body systems, the Lieb–Robinson bound is unavailable, and consequently, the adiabatic continuation operator lacks a simple geometric structure. In these regimes, the conventional SIE theorem ceases to be effective. To overcome this difficulty, we combine the adiabatic theorem [55] with two key tools: the Approximate-Ground-State-Projection formalism [40] and the spectral SIE framework developed in this work. This combination enables us to prove a generalized area law without imposing any assumption of locality.
@@
[Theorem 4 (Informal)] If the target ground state is connected to a trivial ground state that satisfies the generalized area law via an adiabatic path (Assumption 13), then the entanglement entropy across any bipartition is bounded from above by $\bar{{\mathcal{J}}}(V_{AB})/\Delta]^{3}$, where $\Delta$ denotes the spectral gap.
Furthermore, analogous statements can be made regarding low-Schmidt-rank approximations of the ground state with respect to bipartition cuts. To the best of our knowledge, this constitutes the first proof of a generalized area law for ground states in the most general settings without assuming geometric locality.
@@ II.5 Polynomial complexity of the long-range interacting systems
Up to this point, our results have addressed the most general settings, including physical systems without any underlying geometric structure. Nevertheless, the framework we have developed also applies meaningfully in situations where geometric structure is present. Among such cases, a particularly important and nontrivial application concerns the computational complexity of simulating quantum systems with long-range interactions. These systems are typically characterized by interactions that decay with spatial distance $r$ as a power law $r^{-\eta}$ with $\eta>0$, yet a comprehensive understanding of their entanglement properties and simulation complexity remains elusive.
@@
This bound ensures that the entanglement structure generated by long-range interactions remains sufficiently controlled to allow polynomially efficient tensor-network representations. In this work, we treat three fundamental classes of quantum states: ground states, time-evolved states, and thermal equilibrium states. For these states, only quasi-polynomial simulation complexity had previously been established [44, 56, 57]. Our results thus improve the known bounds by showing that, under the above decay condition, these physically relevant states can be captured within polynomial complexity, thereby closing the gap between long-range and short-range interacting systems in one dimension.
@@ II.5.1 Ground states
We begin by considering matrix product state (MPS) approximations of gapped ground states in one-dimensional long-range interacting systems. For such ground states, the entanglement area law has already been established [44, 58], and it is known that the required bond dimension scales as $e^{\log^{5/2}(n)}$, thereby imposing a quasi-polynomial overhead. This limitation arises because most existing approaches rely on constructing an Approximate Ground State Projector (AGSP) using the Chebyshev polynomials of the Hamiltonian [40]. As discussed earlier, polynomial approximations of long-range Hamiltonians typically induce quasi-polynomial complexity, rendering high-precision approximations prohibitively costly. In this work, we overcome this bottleneck by employing time-evolution operators together with the Spectral SIE framework. In particular, we estimate the SE strength of a Gaussian-filter AGSP and demonstrate that such filters yield asymptotically superior approximation properties compared to conventional polynomial-based AGSPs. As a result, we prove that the ground state can be approximated by an MPS with polynomial bond dimension.
@@ II.5.2 Quantum dynamics
In our approach, instead of considering the explicit MPO construction of the time-evolution operator itself, we certify the accuracy of time-evolution simulation using the time-dependent Density Matrix Renormalization Group (t-DMRG) algorithm [59, 60]. This strategy overcomes the expensive task of explicitly constructing the full MPO representation. A key difficulty, previously highlighted by Osborne [32], is that naive error analysis of t-DMRG suffers from exponential error amplification at each time step, undermining reliability (Sec. VIII.4.2). In this work, we address this problem by monitoring the $(1/2)$-Rényi entanglement proxy of the approximated state at each step, thereby guaranteeing controlled error propagation throughout the simulation:
@@ II.5.3 Quantum Gibbs states
A natural question is whether similar techniques can be applied to imaginary time evolution, which is fundamental for studying thermal equilibrium states. This setting, however, presents intrinsic challenges. The key difficulty is that imaginary time evolution does not preserve the norm of the quantum state. Because imaginary-time evolution is non-unitary, even a product-form propagator (e.g., $e^{-\tau H_{A}}\otimes\hat{1}_{B}$) can, after normalization, significantly modify the entanglement spectrum across $A|B$. This illustrates in a striking way why complex-time evolution is qualitatively more difficult than real-time evolution [61, 8]. Consequently, a direct application of the spectral SIE framework only allows efficient MPO (or purified MPS) approximations in high-temperature regimes, where the thermal state remains close to the identity and entanglement is naturally limited.
@@
This result qualitatively improves upon the previously known quasi-polynomial complexity bounds for quantum Gibbs states [56, 57]. On the other hand, the construction of an explicit and efficient algorithm for generating such an MPO remains an open problem.
@@ III.1 General framework
We consider a general many-body qudit system, with the total set of qudits denoted by $\Lambda$. Each qudit is assumed to have a finite Hilbert space dimension. We bipartition the system into two subsystems $A$ and $B$ such that $\Lambda=A\cup B$. For notational simplicity, $A\cup B$ is frequently abbreviated as $AB$, particularly when it appears in a subscript. Additionally, we often use $A_{0},A_{1}$ and $B_{0},B_{1}$ to denote decompositions of $A$ and $B$ into two subsets, such as $A=A_{0}\cup A_{1}$ and $B=B_{0}\cup B_{1}$. For any subset $X\subseteq\Lambda$, we denote the Hilbert space dimension by $\mathcal{D}_{X}$. A product state $|\psi_{A}\rangle\otimes|\psi_{B}\rangle$ is often abbreviated as $|\psi_{A},\psi_{B}\rangle$.
@@
where $H_{A}$ and $H_{B}$ act only on $A$ and $B$, respectively, and $V_{AB}$ represents the interaction term between $A$ and $B$. At this stage, we do not impose any specific locality constraints on the Hamiltonian. In studying the entanglement generation, we typically assume that $H_{A}$ and $H_{B}$ may take arbitrary form (e.g., their norms are unbounded), while only the boundary interaction $V_{AB}$ is restricted (see 1 below).
@@
$E_{\alpha}(\psi)=\frac{1}{1-\alpha}\log\left(\sum_{s}\lambda_{s}^{2\alpha}\right),$ (3)
@@
while $|\Omega\rangle$ denotes the ground state of the Hamiltonian $H$, and $\Delta$ denotes the spectral gap between the ground state and the first excited state. Furthermore, we study the quantum Gibbs state at inverse temperature $\beta$,

### v2

@@ Abstract
How the detailed structure of quantum complexity emerges from quantum dynamics remains a fundamental challenge highlighted by advances in quantum simulators and information processing. As a celebrated milestone, the Small-Incremental-Entangling (SIE) theorem provides a universal constraint on the rate of entanglement generation. While the SIE theorem limits the total amount of entanglement, it leaves a major open problem in fully characterizing the fine entanglement structure. Here we introduce the concept of Spectral-Entangling strength, which captures the structural entangling power of an operator, and use it to establish a new spectral SIE theorem: we derive a universal speed limit for Rényi entanglement growth at $\alpha\geq 1/2$, revealing a robust $1/s^{2}$ decay threshold in the entanglement spectrum. Remarkably, our bound at $\alpha=1/2$ is both qualitatively and quantitatively optimal, establishing the universal threshold beyond which entanglement growth becomes unbounded. This result exposes the detailed structure of Schmidt coefficients and, in turn, enables rigorous truncation-based error control, providing a quantitative link between entanglement structure and computational complexity. Building on these results, our framework establishes a generalized entanglement area law under the adiabatic-path condition, thus extending a central principle of quantum many-body physics to general interactions. As a more practical application, we show that one-dimensional long-range interacting systems admit polynomial bond-dimension approximations for ground states, time-evolved states, and thermal states. This closes the long-standing quasi-polynomial gap and demonstrates that such systems can be simulated with polynomial complexity comparable to short-range models. In particular, by controlling Rényi entanglement, we obtain a rigorous, a priori error guarantee for the time-dependent density-matrix renormalization-group algorithm. Overall, our results extend the scope of the SIE theorem and establish a unified framework that reveals the detailed structure of quantum complexity.
@@ I Introduction
Another well-known information-theoretic constraint that does not rely on geometric locality is the Small-Incremental-Entangling (SIE) theorem [21]. This theorem considers a bipartition of a quantum system into subsystems $A$ and $B$, and asserts that the rate at which entanglement entropy is generated between them can be upper bounded solely in terms of the interaction strength and the Hilbert space dimensions of the directly coupled subsystems $A_{0}\subset A$ and $B_{0}\subset B$. In systems with spatial locality, the SIE bound can be derived from the Lieb–Robinson bound [14]. However, its general validity across arbitrary quantum systems was first conjectured by Kitaev and Bravyi, and later fully proven in 2013 [22, 23, 24, 25]. Because the rate of entanglement growth depends only on the interaction across the boundary between $A$ and $B$, the SIE theorem is often referred to as the Dynamical Area Law. This term emphasizes the fact that entanglement generation in time is governed by the surface—rather than the volume—of the interacting regions. The SIE theorem has served as a key methodological tool in a wide range of fundamental problems, such as supporting the area law conjecture for noncritical ground states in higher-dimensional systems [22, 26], analyzing complexity growth in quantum circuits [27, 28], and characterizing measurement-induced phase transitions [29]. These applications demonstrate the power and universality of the SIE framework. Nevertheless, as we will discuss in the next section, the entanglement generation rate alone is often insufficient to fully capture the complexity of quantum many-body dynamics.
@@
In this work, we move beyond these limitations by establishing a spectral formulation of entanglement growth. Our framework not only provides rigorous dynamical constraints on the entanglement spectrum itself, but also achieves both qualitative and quantitative optimality: at the critical threshold, we establish a universal upper bound corresponding to a $1/s^{2}$ decay of squared Schmidt coefficients, which is sharp and cannot be improved. This perspective sharpens the dynamical area law into a more fine-grained structural principle and extends its reach to broader consequences, including generalized area laws of gapped ground states, efficient matrix-product-state approximations for long-range systems, and rigorous a priori performance guarantees for numerical algorithms such as t-DMRG (time-dependent density-matrix renormalization group). Conceptually, the ability to address the entanglement spectrum directly opens the door to a more refined understanding of quantum complexity, enabling applications that go well beyond entropy-based diagnostics.
@@ II.2 Spectral Small-Incremental-Entangling (SIE)
In general, this decay law cannot be further improved without imposing additional geometric assumptions such as finite-dimensionality or short-range interactions [32, 8, 9]. The result can be derived from an upper bound on the growth rate of the $\alpha$-Rényi entanglement $E_{\alpha}$, as defined precisely in Eq. (3). In order to emphasize the spectral nature of the bound, we refer to this extension as the Spectral Small-Incremental-Entangling (Spectral SIE) theorem.
@@
An important observation is that $\alpha=1/2$ constitutes a universal threshold. For $\alpha<1/2$, the entanglement generation rate becomes unbounded, as proven in Proposition 8. Moreover, for $\alpha=1/2$ the constant $c_{\alpha=1/2}$ is not only universal but also quantitatively optimal, as shown in Proposition 7. A significant advantage of considering the Rényi entanglement with $\alpha<1$ is that it enables rigorous error analysis in terms of Schmidt rank truncation [30, 31], thereby providing a quantitative estimate of how entanglement structure affects the computational complexity of simulating quantum dynamics.
@@ II.3 Low-Schmidt-rank approximation of general operators
As discussed above, it has been established that time-evolved states admit efficient low-rank approximations. A natural question is whether a similar low-rank approximation exists at the level of the time-evolution operator itself. In conclusion, approximating an operator with respect to the operator norm is significantly more challenging. To address this, we demonstrate that there exists a complexity separation between Schmidt-rank approximability of time-evolved quantum states and that of the corresponding time-evolution unitary.
[Proposition 9 (informal)] There exist Hamiltonians $H_{AB}$ with $\bar{{\mathcal{J}}}(V_{AB})=1$ for which the unitary $e^{-iH_{AB}t}$ does not admit any low-rank approximation in a certain (but constant) time regime.
@@ II.4 Generalized entanglement area law
The entanglement area law refers to the property that, for a ground state of a finite-dimensional lattice system partitioned into two regions $A$ and $B$, the entanglement entropy scales with the size of the boundary between the two regions [11, 38]. In particular, the area law conjecture argues that the area law always holds for noncritical ground states with a spectral gap, and it remains one of the central open problems in quantum many-body physics [39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. The notion of an area law can be extended to more general bipartite interaction settings [49]. Specifically, when a system is divided into $A$ and $B$, the Schmidt rank of the interaction operator between $A$ and $B$ can be interpreted as an effective boundary size, thereby allowing a generalized formulation of the area law. For example, in a four-partite setup $(A_{1},A_{0},B_{0},B_{1})$ where the interaction exists only between $A_{0}$ and $B_{0}$, the relevant boundary size is determined by the dimensions of $A_{0}$ and $B_{0}$. Typical examples include generalized area laws for local Hamiltonians defined on graphs, which have been widely discussed in the literature. According to the conventional SIE theorem [22, 23], the generalized area law can be shown to hold for entanglement dynamics even in such general settings. It is also widely known that generalized area laws for mutual information hold for quantum Gibbs states at finite temperature [50, 51]. However, whether the same statement applies to gapped ground states has long been debated, and this question was resolved in the negative by Aharonov et al. in 2014 [49]. At present, aside from special classes [52], it remains a fundamental open problem to determine the precise conditions under which a generalized area law holds.
@@
In this work, we establish that the generalized area law holds under the assumption that an adiabatic path exists [53]. Given a parameter-dependent Hamiltonian $H(s)$, we define an adiabatic path as a continuous trajectory in parameter space that connects a reference Hamiltonian with a known area-law-satisfying ground state to a target Hamiltonian, while maintaining a finite spectral gap throughout. This condition is standard in proofs of the area law for spatially local systems [26, 22, 54], where it is combined with the Lieb–Robinson bound, and where the SIE theorem provides a key estimate on entanglement growth along the adiabatic evolution. In more general settings, such as systems defined on infinite-dimensional graphs or general two-body systems, the Lieb–Robinson bound is unavailable, and consequently, the adiabatic continuation operator lacks a simple geometric structure. In these regimes, the conventional SIE theorem ceases to be effective. To overcome this difficulty, we combine the adiabatic theorem [55] with two key tools: the Approximate-Ground-State-Projection formalism [40] and the spectral SIE framework developed in this work. This combination enables us to prove a generalized area law without imposing any assumption of locality.
@@
[Theorem 4 (Informal)] If the target ground state is connected to a trivial ground state that satisfies the generalized area law via an adiabatic path (Assumption 13), then the entanglement entropy between $A$ and $B$ is bounded from above by $\bar{{\mathcal{J}}}(V_{AB})/\Delta]^{3}$, where $\Delta$ denotes the spectral gap.
Furthermore, analogous statements can be made regarding low-Schmidt-rank approximations of the ground state for the bipartition. To the best of our knowledge, this constitutes the first proof of a generalized area law for ground states in the most general settings without assuming geometric locality.
@@ II.5 Polynomial complexity of the long-range interacting systems
Up to this point, our results have addressed the most general settings, including physical systems without any underlying geometric structure. Nevertheless, the framework we have developed also applies in scenarios where geometric structure is present. Among such cases, a particularly important and nontrivial application concerns the computational complexity of simulating quantum systems with long-range interactions. These systems are typically characterized by interactions that decay with spatial distance $r$ as a power law $r^{-\eta}$ with $\eta>0$, yet a comprehensive understanding of their entanglement properties and simulation complexity remains elusive.
@@
This bound ensures that the entanglement structure generated by long-range interactions is sufficiently constrained to allow polynomially efficient tensor-network representations. In this work, we consider three fundamental classes of quantum states: ground states, time-evolved states, and thermal equilibrium states. For these states, only quasi-polynomial simulation complexity had previously been established [44, 56, 57]. Our results thus improve the known bounds by showing that, under the above decay condition, these physically relevant states can be captured within polynomial complexity, thereby closing the gap between long-range and short-range interacting systems in one dimension.
@@ II.5.1 Ground states
We begin by considering matrix product state (MPS) approximations of gapped ground states in one-dimensional long-range interacting systems. For such ground states, the entanglement area law has already been established [44, 58], and it is known that the required bond dimension scales as $e^{\log^{5/2}(n)}$, thereby imposing a quasi-polynomial overhead. This limitation arises because most existing approaches rely on constructing an Approximate Ground State Projector (AGSP) using the Chebyshev polynomials of the Hamiltonian [40]. As discussed earlier, polynomial approximations of long-range Hamiltonians typically induce quasi-polynomial complexity, rendering high-precision approximations prohibitively costly. In this work, we overcome this bottleneck by employing time-evolution operators together with the spectral SIE framework. In particular, we estimate the SE strength of a Gaussian-filter AGSP and demonstrate that such filters yield asymptotically better approximations than conventional polynomial-based AGSPs. As a result, we prove that the ground state can be approximated by an MPS with polynomial bond dimension.
@@ II.5.2 Quantum dynamics
In our approach, instead of considering the explicit MPO construction of the time-evolution operator itself, we certify the accuracy of time-evolution simulation using the time-dependent Density Matrix Renormalization Group (t-DMRG) algorithm [59, 60]. This strategy overcomes the expensive task of explicitly constructing the full MPO representation. A key difficulty, previously highlighted by Osborne [32], is that naive error analysis of t-DMRG suffers from exponential error amplification at each time step, undermining reliability (Sec. VIII.4.2). In particular, while truncation-based diagnostics (e.g., the discarded weight) provide a posteriori accuracy checks during the simulation [59], a general a priori upper bound on the total error derived solely from the dynamics of a given local Hamiltonian has not been established for standard t-DMRG.
In this work, we address this problem by monitoring the $(1/2)$-Rényi entanglement proxy of the approximated state at each step, thereby guaranteeing controlled error propagation throughout the simulation:
@@ II.5.3 Quantum Gibbs states
A natural question is whether similar techniques apply to imaginary time evolution, which is fundamental for studying thermal equilibrium states. This setting, however, presents intrinsic challenges. The key difficulty is that imaginary time evolution does not preserve the norm of the quantum state. Because imaginary-time evolution is non-unitary, even a product-form propagator (e.g., $e^{-\tau H_{A}}\otimes\hat{1}_{B}$) can, after normalization, significantly modify the entanglement spectrum across $A|B$. This illustrates in a striking way why complex-time evolution is qualitatively more difficult than real-time evolution [61, 8]. Consequently, a direct application of the spectral SIE framework allows efficient MPO (or purified MPS) approximations only in high-temperature regimes, where the thermal state remains close to the identity and entanglement is naturally limited.
@@
This result qualitatively improves upon the previously known quasi-polynomial complexity bounds for quantum Gibbs states [56, 57]. By contrast, designing an explicit and efficient algorithm to generate such an MPO remains open.
@@ III.1 General framework
We consider a general many-body qudit system, with the total set of qudits denoted by $\Lambda$. Each qudit is assumed to have a finite Hilbert space dimension. We bipartition the system into two subsystems $A$ and $B$ such that $\Lambda=A\sqcup B$. For notational simplicity, $A\sqcup B$ is frequently abbreviated as $AB$, particularly when it appears in a subscript. Additionally, we often use $A_{0},A_{1}$ and $B_{0},B_{1}$ to denote decompositions of $A$ and $B$ into two subsets, such as $A=A_{0}\sqcup A_{1}$ and $B=B_{0}\sqcup B_{1}$. For any subset $X\subseteq\Lambda$, we denote the Hilbert space dimension by $\mathcal{D}_{X}$. A product state $|\psi_{A}\rangle\otimes|\psi_{B}\rangle$ is often abbreviated as $|\psi_{A},\psi_{B}\rangle$.
@@
where $H_{A}$ and $H_{B}$ act only on $A$ and $B$, respectively, and $V_{AB}$ represents the interaction term between $A$ and $B$. At this stage, we do not impose any specific locality constraints on the Hamiltonian. In studying the entanglement generation, we typically assume that $H_{A}$ and $H_{B}$ may take arbitrary form (e.g., their norms are unbounded), while only the boundary interaction $V_{AB}$ is restricted (see Def. 1 below).
@@
$E_{\alpha}(\psi)=\frac{1}{1-\alpha}\log\left(\sum_{s}\lambda_{s}^{2\alpha}\right)=\frac{1}{1-\alpha}\log\left[{\rm tr}_{A}\left(\rho^{\alpha}_{A}\right)\right],$ (3)
@@
while $|\Omega\rangle$ denotes the ground state of the Hamiltonian $H$, and $\Delta$ denotes the spectral gap between the ground state and the first excited state. Furthermore, we study the quantum Gibbs state at inverse temperature $\beta$: